# In-ice Askaryan Emission from Air Showers: Implications for Radio Neutrino Detectors

### v1

@@ 1 Introduction
Detection of neutrinos at ultra-high energies (UHE, ) would open a new window to the most violent phenomena in our universe [1, 2, 3]. The detection of neutrinos is challenging, though, due to their small interaction cross-section and the rapidly decreasing neutrino flux towards higher energies. Radio detection remains the most promising technique at these energies, allowing for cost-efficient instrumentation of large volumes [4]. In particular, the in-ice radio detection technique, which instruments ice sheets with arrays of compact radio detector stations, is at an advanced development stage. The technology has been successfully developed and explored with the pilot arrays ARA and ARIANNA at the South Pole and Moore’s Bay in Antarctica [5, 6]. The ongoing construction of the Radio Neutrino Observatory in Greenland (RNO-G) makes the detection of the first UHE neutrino a possibility in the coming years [7], and the planned radio array of IceCube-Gen2 will increase the sensitivity by another factor of 10 [8].
@@
Consequently, a decisive observation of this flux will require an even better understanding of all relevant backgrounds. The physical background at UHE is dominated by secondary particles generated by cosmic-ray interactions in the atmosphere. The impulsive radio-frequency signals generated by these secondaries can mimic those of UHE neutrino interactions in the ice through various channels, and efforts to reduce these backgrounds require a detailed understanding of the signature of each mechanism. While radio emission from two of these channels, mainly geomagnetic emission in the air from opposing deflection of during the air-shower development [10] and stochastic energy deposits in the ice by eV muons have been well studied [11, 12], the third channel has only recently been explored [13, 14, 15]: the in-ice Askaryan emission from the air-shower cores that reach the air-ice boundary. In the following, we focus on this background process. We refer the reader to the review of [4] for an overview and a more detailed description of the other relevant background processes for in-ice radio neutrino detection.
Askaryan emission from air-shower cores has additional modes of producing background that may be more difficult to reject and would require a better understanding of the ice properties and more generally concentrated calibration efforts. Reflection layers in the ice, inhomogeneities of the bulk ice-crystal structure, can reflect signals back up that appear to originate from deep in the ice, as expected from UHE neutrino interactions. Paired with the abundant flux of cosmic rays, the background rate can be substantial. Here, we quantify the expected background rate for the first time and discuss its mitigation strategies.
@@
First, this offers a calibration signal for in-ice radio detector signals, which will allow probing of the trigger efficiency, signal identification, and reconstruction algorithms under realistic conditions. The radio emission will reach the deeper antennas of an in-ice radio detector station on a direct path. This geometry is of limited concern as a background because the signal arrival directions are restricted to a small angular window from close to the surface with only a small overlap to the expected signal arrival directions from neutrino-induced in-ice showers. We will later calculate that the expected detection rates are large, with tens to hundreds of detections per station per year. Once the air-shower core Askaryan emission is understood and modeled accurately, it will be a useful calibration signal to demonstrate that a deep detector system is working. The full analysis pipeline that will be applied to neutrinos of triggering and identifying the signal out of the huge number of noise triggers can be probed. Even the reconstruction of the energy and shower direction can be probed on a statistical basis, as we can predict the expected energy and arrival direction distributions with high accuracy. In addition, for a subset of the events, the in-air radio emission generated by the air shower can be detected by the upward-facing LPDA antennas of the shallow detector component, which will provide an independent measurement of the air-shower direction and an estimate of the energy deposited in the atmosphere [21, 22, 23].
Second, natural ice sheets contain reflection layers (see next section) that act as a mirror and reflect the signal back up so that it appears to originate from below. The reflectivity is small and at most 1%, but because cosmic rays are much more abundant than neutrinos and the cosmic-ray energy spectrum extends to higher energies, the detection rate of reflected shower cores might still be substantial. In this article, we quantify the background detection rate, characterize the background, and discuss mitigation strategies.
@@ 2 Review of Reflection Layers
Typically, reflective layers within the ice sheet are attributed to originate in one of four sources. First, yearly summer snow melt forms crusts with a slight overdensity relative to surrounding snow, which then increasingly embed within the ice sheet. The density contrast is of order 1% for snow melt layers, which have buried to depths beneath the firn. Second, episodic processes such as sulfate depositions resulting from terrestrial volcanism can lead to thin (thicknesses typically much less than one skin depth) but highly conductive layers, often visible during inspection of ice cores, with correspondingly high reflectivity. Third, re-alignment of the crystal orientation fabric, over the course of many years, can also result in an evolution of the dielectric permittivity over macroscopic distances within the ice sheet. Finally, depending on wind patterns, sea salts can be blown inland, resulting in fluctuations in the local or ionic (and therefore conductive) concentrations.
@@
6009 ns | m | -48+/-9 |
9674 ns | m | -47+/-8 |
9792 ns | m | -46+/-8 |
9905 ns | m | -54+/-12 |
The existence of internal reflection-layers in the South Polar ice is corroborated by attenuation measurements [31], which show prominent peaks at depths of and . The bi-static vertical measurements transmitted tone bursts of duration. We use the time-integrated reflected power as a function of depth in the ice from that study to calculate the reflectivity of a given layer using an ice attenuation profile appropriate to the South Pole and the assumption that the reflection coefficient of the reflecting surface at the bottom of the ice sheet has . We use the measured ratio of reflected power at a given depth to the reflected power from the bottom of the ice sheet, and the expected attenuation in power during propagation, to calculate , which for both depths is , consistent with Table 1. We note that the depth estimate between different measurements can have significant systematic uncertainties due to the choice of the density/index-of-refraction profile that is used to convert the measured time delays to depth.
@@ 3 Calculation of in-ice radio emission from shower cores
This is shown schematically in Fig. 2.
The remaining energy that penetrates a few tens of meters into the ice will produce a radio signal that can be observed by in-ice radio antennas. However, the lateral spread of the electromagnetic particles upon reaching the air-ice interface will result in a loss of coherence with respect to a shower that is initiated at this location with energy by a single particle, e.g., a neutrino interaction. As we will later use Askaryan emission models that were developed for neutrino-induced in-ice showers, we need to correct for this loss of coherence. The observed air-shower core signal will be smaller by a factor , giving an effective energy of the air-shower core, , as the amplitude of the Askaryan signal scales linearly with the shower energy. The work of [13] found significant radio emission from in-ice shower cores similar to the emission expected from neutrino-induced in-ice showers. In that work, air shower simulations with CORSIKA 7 were combined with in-ice shower simulations using GEANT-4 to simulate the in-ice radio emission. Although no frequency spectrum was shown, the time-domain pulses are very short (), indicating high-frequency content (). Furthermore, we note that radio neutrino detectors optimize the trigger bandwidth to low frequencies of approx. 80- compared to the full detector bandwidth of up to , which has been shown to increase the sensitivity to neutrinos [18]. Hence, a loss in coherence of shower cores at high frequencies is less relevant when it comes to triggering a radio neutrino detector, and for calculating detection rates, our simplified signal modeling only needs to match the true signals in the 80- frequency band.
@@ 3.1 CORSIKA 8 simulation of the radio emission from air-shower cores
The recent development of the CORSIKA 8 code [39, 19, 20] allows us to conveniently simulate the in-ice radio emission from air-shower cores in a single framework. We simulated several vertical air showers with proton primaries of various energies, ranging from to . We approximate the ice by a homogeneous medium with an index-of-refraction of 1.78, without attenuation. The observer positions are placed at a depth of and with horizontal displacements from the shower axis to map relevant viewing angles around the Cherenkov cone.
In Fig. 3, we show the in-ice radio emission generated by a vertical air shower initiated by a proton at the South Pole. In this example, = eV is deposited into the ice. The time-domain signals are filtered to the trigger bandwidth and shown for three different viewing angles. In addition, we show the prediction of the radio emission of a neutrino-induced in-ice NC shower with a shower energy of using the parameterization of [40, 41] which is available through NuRadioMC [42]. We find that the time-domain waveforms are very similar in shape. As expected, in this band, the amplitudes from the proper CORSIKA 8 calculation are smaller with a -factor of 0.4 on the Cherekov cone, which grows to at off-cone. We find that the frequency cutoff is at higher frequencies for air-shower cores, likely due to the larger lateral spread of the shower. As a consequence, the Cherenkov cone is broader, and the -factor increases with the viewing angle. Further, the frequency spectra of Fig. 3 show that the emission from shower cores can even exceed the emission from neutrino-induced showers of the same energy at larger frequencies and off-cone angles.
@@
By analyzing air-showers of different energies, we find that the -factor is most closely tied to the ratio /, which results in an implicit scaling with energy and an inverse scaling with zenith angle as this moves further away from the ground, which reduces . We also observe that the f-factor is generally larger for viewing angles further from the Cherenkov cone. For a vertical proton-induced air shower, we find -factors between 0.2 to 0.7, and for a 20∘ air shower, we find -factors between 0.3 and 0.6. While we take into account that less energy is deposited into the ice for inclined air showers through Eq. 2, we expect the -factor to decrease with larger zenith angles because the shower reaches the ice at a later shower age with increased lateral spread.
@@
While more work needs to be done to properly parameterize the radio emission from air-shower cores, the results from [13, 15] and the new CORSIKA 8 results presented here, indicate that average -factors of around 0.1–0.7 are plausible. In future work, we plan to incorporate the effects of an inhomogeneous ice sheet where the index-of-refraction transitions from approx. 1.4 at the surface to 1.78 at deeper depth into CORSIKA 8 to make more accurate predictions. Then, the emission can be parameterized similar to what has been done for neutrino-induced showers, e.g., [43], which can then directly be used in NuRadioMC to simulate the effective area to air-shower cores.
In the following, we treat as a free parameter. We note that the use of the scaling factor, , is simplistic with respect to a true situation wherein the amount of decoherence depends on several parameters, such as the shower age at the surface, or, more directly, with energy, and the zenith angle. However, an effective- can always be found, averaged over the relevant phase space.
@@ 4.1 Calculation of deposited energy spectrum
We perform the analysis in small bins of cosmic-ray energy of 0.1 in and cosine zenith angle of 0.05. For each cosmic-ray energy bin, we draw 200 realizations from a normal distribution with mean and sigma according to the measured values and calculate the deposited energy using Eq. 2. Additionally, the statistical measurement uncertainties of these quantities are incorporated by randomly shifting the central values of the energy spectrum and distributions within the reported errors. Where necessary, the measured values are extrapolated to cover the relevant phase space. This yields a cosmic-ray energy spectrum as a function of deposited energy, as shown in Fig. 4. At South Pole altitude, the deposited-energy spectrum is offset from that of cosmic rays by about 0.5 in . Fig. 5 shows the deposited-energy spectrum for different cosine zenith bands. Each band covers the same solid angle in the sky. As expected, the flux is highest for vertical showers and decreases for more inclined showers as is further away from the ground and less energy reaches the ground (the atmospheric depth approximately increases with ). Simultaneously, the projected area decreased with . Hence, vertical showers contribute the most to the flux, for a fixed value of , by up to two orders of magnitude compared to one at .
@@ 4.2 Calculation of effective area to shower cores
With the NuRadioMC simulation code [42], we can simulate the effective area to in-ice showers, in particular, the showers initiated by the air shower cores beneath the ice surface. We simulate the exact detector as foreseen for IceCube-Gen2 with the same simulation settings as in previous studies [8, 16]. The IceCube-Gen2 radio array will consist of a hybrid array of shallow components with LPDA antennas installed a few meters below the surface and deep components with antennas installed in narrow holes installed down to a depth of . Both components trigger independently. The shallow component uses a high/low threshold trigger with a local coincidence requirement of two out of the four LPDAs fulfilling the trigger condition within . The deep component is triggered by an interferometric phased array comprising four dipole antennas placed vertically above each other with separation [17]. The trigger thresholds are adjusted to yield a trigger rate of on unavoidable thermal noise fluctuations. The exact simulation settings and NuRadioMC steering files are available at [47]. In Sec. 4.4, we will also show results for a phased array at the site of RNO-G. These will use the same frequency band and thresholds but with a depth of and with the commensurate refractive index profile.
The simulations are performed for reflection layers at , , and . We use the following scheme to directly evaluate all possible reflectivies: In the simulation of the effective area, the reflectivity of these layers was set to unity. Since the reflectivity reduces the signal amplitude linearly, we can infer the rates over a range of reflectivities by simply rescaling the simulated energies.111For example, a eV shower that triggers a station via a reflection off of a 0 dB reflector would give an equivalent signal as a eV shower with a -40 dB reflector. As the signal amplitude also depends linearly on the shower energy, we can incorporate any reflectivity later by reducing the effective shower energy accordingly.

### v2

@@ 1 Introduction
Detection of neutrinos at ultra-high energies (UHE, ) would open a new window to the most violent phenomena in our universe [1, 2, 3]. The detection of neutrinos is challenging, though, due to their small interaction cross-section and the rapidly decreasing neutrino flux towards higher energies. Radio detection remains the most promising technique at these energies, allowing for cost-efficient instrumentation of large volumes [4]. In particular, the in-ice radio detection technique, which instruments ice sheets with arrays of compact radio detector stations, is at an advanced development stage. The technology has been successfully developed and explored with the pilot arrays RICE, ARA and ARIANNA at the South Pole and Moore’s Bay in Antarctica [5, 6, 7]. The ongoing construction of the Radio Neutrino Observatory in Greenland (RNO-G) makes the detection of the first UHE neutrino a possibility in the coming years [8], and the planned radio array of IceCube-Gen2 will increase the sensitivity by another factor of 10 [9].
@@
Consequently, a decisive observation of this flux will require a precise understanding of all relevant backgrounds. The physical background at UHE is dominated by secondary particles generated by cosmic-ray interactions in the atmosphere. The impulsive radio-frequency signals generated by these secondaries can mimic those of UHE neutrino interactions in the ice through various channels, and efforts to reduce these backgrounds require a detailed understanding of the signature of each mechanism. While radio emission from two of these channels, in-air radio emission which originates mainly from geomagnetic emission from opposing deflection of during the air-shower development [11, 12, 13] and stochastic energy deposits in the ice by eV muons generating Askaryan emission have been well studied [14, 15], the third channel has only recently been explored [16, 17, 18]: the in-ice Askaryan emission from the air-shower cores that reach the air-ice boundary [19, 20]. In the following, we focus on this background process. We refer the reader to the review of [4] for an overview and a more detailed description of the other relevant background processes for in-ice radio neutrino detection.
Askaryan emission from air-shower cores has additional modes of producing background that may be more difficult to reject and would require a better understanding of the ice properties and more generally concentrated calibration efforts compared to the other background channels. Reflection layers in the ice, inhomogeneities of the bulk ice-crystal structure, can reflect signals back up that appear to originate from deep in the ice, as expected from UHE neutrino interactions. Paired with the abundant flux of cosmic rays, the background rate can be substantial. Here, we quantify the expected background rate for the first time and discuss its mitigation strategies.
@@
First, this offers a calibration signal for in-ice radio detector signals, which will allow probing of the trigger efficiency, signal identification, and reconstruction algorithms under realistic conditions. The radio emission will reach the deeper antennas of an in-ice radio detector station on a direct path. This geometry is of limited concern as a background because the signal arrival directions are restricted to coming from above with zenith angles in a given range with small overlap to the expected signal arrival directions from neutrino-induced in-ice showers. We will later calculate that the expected detection rates are large, with tens to hundreds of detections per station per year. Once the air-shower core Askaryan emission is understood and modeled accurately, it will be a useful calibration signal to demonstrate that a deep detector system is working. The full analysis pipeline responsible for triggering on neutrinos and distinguishing genuine signals from the overwhelming background noise can be systematically validated. Even the reconstruction of the energy and shower direction can be probed on a statistical basis, as we can predict the expected energy and arrival direction distributions with high accuracy. In addition, for a subset of the events, the in-air radio emission generated by the air shower can be detected by the upward-facing LPDA antennas of the shallow detector component, which will provide an independent measurement of the air-shower direction and an estimate of the energy deposited in the atmosphere [26, 27, 28].
Second, natural ice sheets contain reflection layers (see next section) that act as a mirror and reflect the signal back up so that it appears to originate from below. The reflectivity is small and at most 1%, but because cosmic rays are much more abundant than neutrinos, have a much larger interaction cross-section, and the cosmic-ray energy spectrum extends to high energies, the detection rate of reflected shower cores might still be substantial. In this article, we quantify the background detection rate, characterize the background, and discuss mitigation strategies.
@@ 2 Review of Reflection Layers
Typically, reflective layers within the ice sheet are attributed to originate in one of four sources. First, yearly summer snow melt forms crusts with a slight overdensity relative to surrounding snow, which then increasingly embed within the ice sheet. The density contrast is typically not more than 1% for snow melt layers, which then embed within, and below the firn. Such yearly crusts are unlikely to produce significant reflections. Second, episodic processes such as sulfate depositions resulting from terrestrial volcanism can lead to thin (thicknesses typically less than one skin depth) but highly conductive layers. These are often visible during inspection of ice cores, with correspondingly high reflectivity and are perhaps most likely responsible for the internal echoes observed in radar surveys. Third, re-alignment of the crystal orientation fabric, over the course of many years, can also result in an evolution of the dielectric permittivity over macroscopic distances within the ice sheet. Finally, depending on wind patterns, sea salts can be blown inland, resulting in fluctuations in the local or ionic (and therefore conductive) concentrations.
@@
| 6009 ns | m | -48+/-9 |
| 9674 ns | m | -47+/-8 |
| 9792 ns | m | -46+/-8 |
| 9905 ns | m | -54+/-12 |
In another study, bi-static vertically-penetrating radio pulses at were transmitted into South Polar ice as tone bursts of duration [36]. The attenuation length of the ice was determined from time-integrated signals that were reflected from the ice-rock interface at the bottom of the ice sheet. However, the reflected power was collected as a function of depth throughout the ice sheet. These data show prominent peaks at depths of and , corroborating the existence of internal reflection layers. We use the time-integrated reflected power as a function of depth in the ice from that study to calculate the reflectivity of a given layer using an ice attenuation profile appropriate to the South Pole and the assumption that the reflection coefficient of the reflecting surface at the bottom of the ice sheet has . We use the measured ratio of reflected power at a given depth to the reflected power from the bottom of the ice sheet, and the expected attenuation in power during propagation, to calculate , which for both depths is , consistent with Table 1. We note that the depth estimate between different measurements can have significant systematic uncertainties due to the choice of the density/index-of-refraction profile that is used to convert the measured time delays to depth.
@@ 3 Calculation of in-ice radio emission from shower cores
where is a normalization constant given by the integral of eq. 1 over all .
This function is shown schematically in the lower panel of Fig. 2.
The remaining energy that penetrates a few tens of meters into the ice will produce a radio signal that can be observed by in-ice radio antennas. However, the lateral spread of the electromagnetic particles upon reaching the air-ice interface will result in a loss of coherence with respect to a shower that is initiated at this location with energy by a single particle, e.g., a neutrino interaction. As we will later use Askaryan emission models that were developed for neutrino-induced in-ice showers, we need to correct for this loss of coherence. The observed air-shower core signal will be smaller by a factor , giving an effective energy of the air-shower core, , as the amplitude of the Askaryan signal scales linearly with the shower energy. The work of [16] found significant radio emission from in-ice shower cores similar to the emission expected from neutrino-induced in-ice showers. In that work, air shower simulations with CORSIKA 7 were combined with in-ice shower simulations using GEANT-4 to simulate the in-ice radio emission, which was named FAERIE simulations by the authors. Although no frequency spectrum was shown, the time-domain pulses are very short (), indicating high-frequency content (). Furthermore, we note that radio neutrino detectors optimize the trigger bandwidth to low frequencies of approx. 80- compared to the full detector bandwidth of up to , which has been shown to increase the sensitivity to neutrinos [23]. Hence, a loss in coherence of shower cores at high frequencies is less relevant when it comes to triggering a radio neutrino detector, and for calculating detection rates, our simplified signal modeling only needs to match the true signals in the 80- frequency band.
@@ 3.1 CORSIKA 8 simulation of the radio emission from air-shower cores
The recent development of the CORSIKA 8 code [24, 25, 44] allows us to conveniently simulate the in-ice radio emission from air-shower cores in a single framework (in contrast to FAERIE). We simulated several vertical air showers with proton primaries of various energies, ranging from to . We approximate the ice by a homogeneous medium with an index-of-refraction of 1.78, without attenuation. The observer positions are placed at a depth of and with horizontal displacements from the shower axis to map relevant viewing angles around the approximate Cherenkov cone with a spacing. We only calculate the radio emission from in-ice shower tracks, i.e., we don’t show and consider the in-air radio emission in this work.
In Fig. 3, we show the in-ice radio emission generated by a vertical air shower initiated by a proton at the South Pole. In this example, = eV is deposited into the ice. The time-domain signals are filtered to the trigger bandwidth and shown for three different viewing angles. The right panel shows the observer with the largest signal strength, i.e., the observer which is closest to the Cherenkov cone. The center and left panels show the emission at two selected directions at and away from the observer with the largest signal strength.
We note that our approach, which considers only the in-ice portion of the particle tracks, in principle introduces additional contributions from the sudden appearance of charges–a contribution to transition radiation (see e.g. [48]). However, our results do not show any indications of such effects. Instead, the observed emission is consistent with expectations from Cherenkov radiation: the amplitude increases with frequency up to a cut-off frequency that peaks at the Cherenkov angle and rapidly decreases away from it. This supports the validity of our current approach. Nonetheless, it would be interesting to investigate potential contributions from transition radiation in future studies.
In Fig. 3, we also present predictions for the radio emission from a neutrino-induced in-ice NC shower with a shower energy of , using the parameterization from [45, 46], commonly referred to as Alvarez2009 and available through NuRadioMC [47]. The comparison, however, is complicated by the fact that the Cherenkov angle is not well-defined for nearby showers. In contrast, neutrino-induced in-ice showers typically occur at distances on the order of (), where they can be reasonably approximated as point-like emitters [47]. To achieve the most accurate comparison, the Alvarez2009 parameterization is evaluated at the Cherenkov angle for the rightmost panel of Fig. 3, and at and from the Cherenkov angle for the other two panels. Since our CORSIKA 8 simulations placed observers at intervals along the viewing angle, the observer with the maximum signal strength can be at most away from the Cherenkov cone.
Within the narrow trigger band, we find that the time-domain waveforms are very similar in shape. As expected, the amplitudes from the shower cores are reduced compared to the neutrino-induced in-ice showers, with an -factor of 0.4 near the Cherenkov cone. This factor increases to at off-cone and reaches at off-cone. We also observe that the cutoff frequency at the approximate Cherenkov angle for shower cores occurs at lower frequencies than for in-ice showers. However, the decline in cutoff frequency with increasing off-cone angle is slower for shower cores, leading to an inversion at larger off-cone angles. Consequently, the Cherenkov cone is broader, and the -factor increases with the viewing angle.
@@
By analyzing air showers of different energies, we find that the -factor is most closely tied to the shower age at the surface, which we quantify by the ratio / which captures both the zenith angle and dependency. We observe that the f-factor is generally larger for viewing angles further from the Cherenkov cone. For a vertical proton-induced air shower, we find -factors between 0.3 to 0.7, and for a 20∘ air shower, we find -factors between 0.2 and 0.7. While we take into account that less energy is deposited into the ice for inclined air showers through Eq. 2, we expect the -factor to decrease with larger zenith angles because the shower reaches the ice at a later shower age with increased lateral spread.
@@
While more work needs to be done to properly parameterize the radio emission from air-shower cores, the results from the FAERIE simulations [16, 18] and the new CORSIKA 8 results presented here, indicate that average -factors of around 0.1–0.9 are plausible. In future work, we plan to incorporate the effects of an inhomogeneous ice sheet where the index-of-refraction transitions from approx. 1.4 at the surface to 1.78 at deeper depth into CORSIKA 8 to make more accurate predictions. Then, the emission can be parameterized similar to what has been done for neutrino-induced showers, e.g., [49], which can then directly be used in NuRadioMC to simulate the effective area to air-shower cores. We note that using CORSIKA 8 simulations directly to calculate the effective area is not feasible because of the long runtimes. A single CORSIKA 8 shower simulation with the calculation of radio emission can take up to a month, but the calculation of the effective area with sufficient statistical precision requires the simulation of hundreds of thousands of showers, particularly at low energies where the trigger probability is small (see Sec. 4.2).
In the following, we present all results as a function of the -factor, enabling detection rates to be determined for any given -value. We recognize that using a single scaling factor, , is a simplification, as the degree of decoherence depends on multiple parameters, including shower age at the surface, energy, , and zenith angle. Nevertheless, an effective -factor can be defined by averaging over the relevant phase space. When a specific -value is required for presenting results, we adopt , which is our best estimate for the effective value. The key finding from the study above is that the uncertainty introduced by the -factor approximation leads to variations by a factor of 2–3, but not by orders of magnitude.
@@ 4.1 Calculation of deposited energy spectrum
We perform the analysis in small bins of cosmic-ray energy of 0.1 in and cosine zenith angle of 0.05. For each cosmic-ray energy bin, we draw 200 realizations from a normal distribution with mean and sigma according to the measured values and calculate the deposited energy using Eq. 2. Additionally, the statistical measurement uncertainties of these quantities are incorporated by randomly shifting the central values of the energy spectrum and distributions within the reported errors. Where necessary, the measured values are extrapolated to cover the relevant phase space. This yields a cosmic-ray energy spectrum as a function of deposited energy, as shown in Fig. 4. At South Pole altitude, the deposited-energy spectrum is offset from that of cosmic rays by about 0.5 in . Fig. 5 shows the deposited-energy spectrum for different cosine zenith bands. Each band covers the same solid angle in the sky. As expected, the flux is highest for vertical showers and decreases for more inclined showers as is further away from the ground and less energy reaches the ground (the atmospheric depth approximately increases with ). Hence, vertical showers contribute the most to the flux, for a fixed value of , by over three orders of magnitude compared to one at .
@@ 4.2 Calculation of effective area to shower cores
With the NuRadioMC simulation code [47], we can simulate the effective area to in-ice showers, in particular, the showers initiated by the air shower cores beneath the ice surface. We simulate the exact detector as foreseen for IceCube-Gen2 with the same simulation settings as in previous studies [9, 21]. The IceCube-Gen2 radio array will consist of a hybrid array of shallow components with LPDA antennas installed a few meters below the surface and deep components with antennas installed in narrow holes installed down to a depth of . Both components trigger independently. The shallow component requires a high and a low threshold crossing within with a local coincidence requirement of two out of the four LPDAs fulfilling the trigger condition within . The deep component is triggered by an interferometric phased array comprising four dipole antennas placed vertically above each other with separation [22]. Multiple beams covering the relevant zenith angle range are formed by coherently summing the signal from the four antennas to produce synthetic waveforms with increased signal-to-noise ratio. The trigger condition is fulfilled if the integrated power in a sliding window for any of the synthetic waveforms crosses a certain threshold. The trigger thresholds are adjusted to yield a trigger rate of on unavoidable thermal noise fluctuations. In the following, we refer to these two station components as "LPDA 2of4 100Hz" and "PA 4ch 100Hz". The exact simulation settings and NuRadioMC steering files are available at [53]. In Sec. 4.4, we will also show results for a phased array at the site of RNO-G. These will use the same frequency band and thresholds but with a depth of and with the commensurate refractive index profile.
We calculate the Askaryan emission using the frequency domain parameterization of in-ice showers Alvarez2009 [45] as implemented in NuRadioMC [47]. We will later correct the effective areas by the correction factor (see next subsection). We describe the index-of-refraction depth profile, i.e., the transition from snow to solid ice in the upper (), with an exponential model and calculate the signal trajectories through ray tracing [47] as done in previous studies, e.g., [9].
The simulations are performed for reflection layers at , , and . We use the following scheme to directly evaluate all possible reflectivities: In the simulation of the effective area, the reflectivity of these layers was set to unity. Since the reflectivity reduces the signal amplitude linearly, we can infer the rates over a range of reflectivities by simply rescaling the simulated energies.111For example, a eV shower that triggers a station via a reflection off of a 0 dB reflector would give an equivalent signal as a eV shower with a -40 dB reflector. As the signal amplitude also depends linearly on the shower energy, we can incorporate any reflectivity later by reducing the effective shower energy accordingly.